### v1

# Quantifying Preferences of Vision-Language Models via Value Decomposition in Social Media Contexts

Abstract
The rapid advancement of Vision-Language Models (VLMs) has expanded multimodal applications, yet evaluations often focus on basic tasks like object recognition, overlooking abstract aspects such as personalities and values. To address this gap, we introduce Value-Spectrum, a visual question-answering benchmark aimed at assessing VLMs based on Schwartz’s value dimensions that capture core values guiding people’s beliefs and actions across cultures. We constructed a vectorized database of over 50,000 short videos sourced from TikTok, YouTube Shorts, and Instagram Reels, covering multiple months and a wide array of topics such as family, health, hobbies, society, and technology. We also developed a VLM agent pipeline to automate video browsing and analysis. Benchmarking representative VLMs on Value-Spectrum reveals significant differences in their responses to value-oriented content, with most models exhibiting a preference for hedonistic topics. Beyond identifying natural preferences, we also explored the ability of VLM agents to adopt specific personas when explicitly prompted, revealing insights into the models’ adaptability in role-playing scenarios. These findings highlight the potential of Value-Spectrum as a comprehensive evaluation set for tracking VLM advancements in value-based tasks and for developing more sophisticated role-playing AI agents.
Quantifying Preferences of Vision-Language Models via Value Decomposition in Social Media Contexts
Jingxuan Li1,*, Yuning Yang1,*, Shengqi Yang2, Yizhou Zhao1, Ying Nian Wu1, 1University of California, Los Angeles, 2Los Alamos National Laboratory, Correspondence: madili@g.ucla.edu
11footnotetext: The authors contributed equally to this work and are listed in alphabetical order by first name.
1 Introduction
Figure 1: Exploring Value-Driven Role-Playing in Vision-Language Models. This study investigates how VLMs adopt assigned personas to align value traits and preferences within the social media contexts.
Figure 2: Overview of Video Distribution of Value-Spectrum Dataset. We collected an abundance of short videos relevant to 10 Schwartz values. The area of circles centered at each keyword represents the number of relevant videos in the database.
Vision-Language Models (VLMs), which align Large Language Models (LLMs) with pre-trained vision encoders through cross-modal alignment training, have shown impressive perceptual and cognitive capabilities in tasks like visual question answering and image captioning, especially when processing everyday images.  Zhou et al. (2019); Radford et al. (2021); Zhang et al. (2023); Lu et al. (2024) As these abilities advance, VLMs are being deployed globally, bringing to light concerns within the responsible AI community about the potential misalignment of synthetic preferences Li et al. (2024), personalities Serapio-García et al. (2023), and values Ren et al. (2024).
Building on these capabilities, questions arise about whether VLMs, like LLMs, exhibit inherent personality traits or synthetic preferences that align with human-designed values. Addressing this gap, our study explores two key questions: (1) Do VLMs exhibit distinct preference traits? (2) Can we embed a human-designed persona into VLMs to influence their behavior and preferences? To systematically evaluate these questions, we propose a framework that evaluates VLM personality traits through an analysis of their value systems. In human psychology, values are guiding principles that influence attitudes, beliefs, and traits, forming a core component of personality Schwartz (2012). By evaluating how VLMs prioritize these values, we can gain insights into their preference traits and alignment with human-designed personas.
We introduce Value-Spectrum, a benchmark designed to reveal value-related traits in VLMs through value-based keyword searches. Using VLM agents embedded within social media platforms, we compiled a dataset of $50,191$ unique video records from different social media platforms, each segment aligned with distinct value dimensions from the theory of basic human values (see Fig. 2). By presenting the screenshots of these videos to various VLM agents and asking value-based preference questions, we quantified preference differences between models.
Figure 3: VLM Agent for Social Media Video Collection and Interaction. Our pipeline evaluates and stores VLM’s value preferences across social media platforms. This pipeline uses a VLM agent to collect and process videos from social media platforms like Instagram, TikTok, and YouTube. AI models such as GPT, Gemini, Qwen, and Claude interpret the content of the videos. The system then presents a video to the user and asks if the content is interesting. Based on the user’s yes/no response, the agent either continues watching the video or moves to the next one. This creates an interactive system that tailors the video collection experience to user preferences.
Building upon this benchmark, we systematically evaluated several state-of-the-art VLMs to analyze their alignment with value dimensions. Our findings revealed a shared tendency across models: most exhibited a strong inclination towards certain values, such as Hedonism, with consistently positive responses up to 75% of the time. However, alignment varied significantly across models. For instance, CogVLM and Qwen-VL-Plus demonstrated broad and consistent alignment across all value dimensions, while closed-source models like GPT-4o and Claude-3.5 showed distinct preferences, favoring values such as Self-direction and Hedonism. In contrast, Blip-2 exhibited minimal alignment across all value dimensions, highlighting its limitations in expressing stable preferences. These findings highlight diverse capabilities among VLMs in aligning with human value dimensions but also reveal challenges in inducing consistent, human-designed personas across models.
To address these challenges, we propose two strategies—Simple and ISQ—to systematically adjust value decompositions and induce specific personas within VLMs. By evaluating the effectiveness of these strategies across multiple platforms, our experiments show that TikTok serves as an optimal testing environment for inducing VLM personalities, with models demonstrating the strongest alignment under ISQ. Notably, Claude-3.5 achieved the largest gains with ISQ, whereas Blip-2 showed no improvement under either strategy, underscoring fundamental differences in model adaptability. These results validate our approach and provide a pathway for optimizing VLMs’ personality alignment in real-world applications.
In brief, this work makes the following contributions:
  • We present a dataset of over $50$k short videos spanning diverse topics, social media platforms, and release dates, designed to systematically evaluate the personalities and preferences of VLMs.
  • We propose Value-Spectrum, a benchmark for quantifying VLM value preferences, using social media-based assessments to reveal stable traits across different VLM architectures.
  • We embed specific role-play preferences into VLMs using two strategies(simple and ISQ) to adjust value decompositions, achieving improved personality alignment in cross-platform and real-world tests.
2 Related work
2.1 Vision-Langauge Agents
A Vision-Language Model (VLM) is a fusion of vision and natural language models. Its inputs are images and textual descriptions and it learns to discover the knowledge from the two modalities. The recent development of large VLMs is rapidly advancing the field of AI. These models have the potential to revolutionize various industries and tasks, showcasing their power in plot and table identifying (Liu et al., 2022), visual-question answer (VQA) (Hu et al., 2024), image captioning (Bianco et al., 2023), and e.t.c. Following Niu et al. (2024), the environment for VLM agents to interact with social media can be constructed, we design an automated control pipeline that guides the agent to continuously interact with social networks.
2.2 Computational Social Science
The intersection of social media and computational social science has emerged as a dynamic field of research (Chen et al., 2023). Dialogues and social interactions, with their vast user base and intricate networks of connections, offer a large database to study human behaviors (Christakis and Fowler, 2013), social relationships (Qiu et al., 2021), and social networks (Zhang and Amini, 2023). Researchers in computational social science apply advanced computational techniques, such as machine learning, natural language processing, and network analysis, to analyze massive datasets extracted from social media platforms. These analyses provide insights into various phenomena, including information diffusion (Jiang et al., 2014), opinion formation (Xiong and Liu, 2014), and collective behavior (Pinheiro et al., 2016).
2.3 Sentiment, Personality, and Value
The community has been using machine learning-based models to study human sentiment (Malviya et al., 2020), personality (Stachl et al., 2020), and value (Qiu et al., 2022). revious studies focused on human personality classification (e.g. Myers-Briggs Type Indicator (MBTI)) and machine behaviors (i.e., LLMs’ personality Serapio-García et al. (2023)). Inspired by recent studies on indirectly revealing AI agent’s personalities by physcogical exams Jiang et al. (2024), Questionnaires Huang et al. (2023), and cultural perspectives Kovač et al. (2023), we use the new perspective of revealing VLM’s persona by examining machine behaviors and personalities to evaluate their performance on mainstream social media platforms.
Figure 4: Schwartz Value-Based Video Retrieval Pipeline. This pipeline retrieves videos based on Schwartz values by associating each value with relevant words, such as linking Hedonism to topics like vacation. These word associations are converted into vector queries, retrieving matching video content from a database.
3 Data Collection
Inspired by ScreenAgent Niu et al. (2024), our work leverages a VLM-driven graphical user interface (GUI) agent to autonomously navigate popular social media platforms. This agent conducts random walks through social media platforms where it observes and captures video links alongside screenshots. The data collected are stored in a vector database Han et al. (2023), creating a structured repository optimized for value decomposition and efficient retrieval. We aim to analyze VLM behavior across diverse social contexts and reveal VLMs’ preferences. The automated data collection (see Fig 3) process efficiently fetched a large volume of diverse content, enabling the scope and depth of the analysis that traditional manual collection methods could not achieve.
The resulting dataset comprises $50,191$ video links sourced from Instagram ($32\%$), YouTube ($29\%$), and TikTok ($39\%$). Each entry includes the video link, a representative screenshot, and meta-information such as platform name and post date, capturing a comprehensive snapshot of content posted between July 31st, 2024, and October 31st, 2024. By distributing data evenly across these platforms, we enable balanced analysis and facilitate unbiased value decomposition across social media content. This innovative data set empowers researchers to explore the behavior of VLM in a systematic and organized way, fostering deeper insights into model interpretation and the dynamics of social media.
To examine the distribution of video themes in this data set, we take the screenshot of each video in the beginning as the representation of the video content. This selection is justified in Section 6. We then vectorize the image and define its relevance to a specific keyword as the L2 distance. In Figure 2 we present the abundances of videos that are relevant to ten Schwartz Values. Specifically, for each Schwartz Value we populate representative keywords generated from GPT-4o. The area of the transparent circle is proportional to the number of videos that lie within a distance of $1.5$ to the corresponding normalized keyword vector. Through this simple diagram we find that videos relevant to these Schwartz Value Dimensions Achievement, Hedonism, and Power appear most frequently, while videos about Tradition are relatively rare.
4 Evaluating VLM’s Preferences
Extending the idea of analysing Large Langauge Models’ human likeness (Shanahan et al., 2023; Kovač et al., 2023) to Vision-Language Models (VLMs) with both pictorial and textual inputs, we asks: Do VLMs also exhibit inherent preferences?
To answer this question, We explore a diverse set of VLMs including GPT-4o OpenAI (2023) , Gemini-1.5 Team et al. (2023), Claude-3.5 Anthropic (2023), Qwen-VL-Plus Bai et al. (2023), CogVLM Wang et al. (2023), and Blip-2 Li et al. (2023) to assess their value preferences. We quantified VLMs’ preferences by evaluating their attitudes toward the $10$ Schwartz values: Hedonism, Self-direction, Security, Achievement, Stimulation, Tradition, Benevolence, Universalism, Conformity, and Power. This approach enables us to construct a comprehensive profile of each model’s value preferences and to identify its unique value traits.
After constructing a vector database (as described in Section 3) to retrieve videos based on specific Schwartz values, we analyzed and compared the responses and attitudes of each model toward video screenshots corresponding to those values. Our analysis revealed the extent to which each value captures the VLMs’ attention, uncovering both similarities and differences across models and highlighting distinct inclinations and sentiments within each VLM.
4.1 Preference Retrieval
To evaluate a VLM’s preference for a specific Schwartz value, we collected each model’s responses to images associated with several keywords related to the value (see Figure 4). For instance, we selected the keywords Equality, Globe, and Handshake for the Universalism dimension because they closely align with its core principles of fairness and global awareness. For each keyword linked to the value, each model reviews five images and answers their attitude towards each image.
We retrieve the preference score of each VLM on the given pictorial input according to the context (You are scrolling on video shorts and see this screenshot) and the following prompts: (1) Is this content interesting? (2) What is the thing that makes this picture interesting? (3) Describe this image in English briefly.
The answer to the first question is processed into either yes ($1$) or no (0), and the average score is calculated to evaluate the intensity of the model’s preference for a given value (e.g., Universalism), ranging from 00 to $1$.
4.2 Preference Patterns
We evaluated and visualized the preference dimensions, identifying three distinct patterns:
Figure 5: Each VLM’s preference scores towards the 10 Schwartz values. The scores range from 0 to 1, with higher scores indicating stronger preferences for the corresponding values.
(1) Global Pattern: After summarizing the preference score across all VLMs, we found most of them tend to prioritize certain values, such as Hedonism, over others. The results indicate a general preference for Hedonism and Self-direction, while showing less excitement for Conformity and Power. This trend reflects current societal tendencies toward the pursuit of pleasure and personal goal in life Wrosch and Scheier (2003). The specific ranking is presented in Figure 6.
(2) Range Consistency: As shown in Figure 5, each model’s preference scores remain within a narrow band of approximately ±0.1 around a central value. Models display varying levels of engagement with the content: some, like Blip-2, are more reserved, occupying a smaller area on the plot with lower preference scores, while others, like CogVLM, show greater enthusiasm, demonstrating interest across all inputs with higher preference scores.
(3) Individual Model Variations: When analyzed individually, some models, such as Qwen VL Plus, exhibit consistently high preferences across all 10 Schwartz values. In contrast, other models, like Claude, displays more specific and nuanced preferences as indicated in the standard deviation of responses given by the model across all values (Figure 7).
Value GPT4o Q-wen vl plus Blip 2 CogVLM Gemini 1.5 Claude 3.5
Self-direction 96±8 99±1 28±20 90±13 56±23 74±22
Universalism 90±13 94±9 18±17 98±6 48±18 62±23
Benevolence 86±24 98±6 20±20 99±1 48±16 72±29
Stimulation 90±13 98±6 26±22 98±6 54±27 70±18
Power 84±15 88±13 28±10 92±13 42±26 56±23
Achievement 75±31 96±12 22±19 99±1 66±16 78±21
Hedonism 92±10 98±6 10±13 92±10 64±17 86±16
Conformity 78±28 96±8 14±18 94±9 50±22 66±27
Tradition 84±12 96±8 22±17 96±8 56±28 64±33
Security 90±16 98±6 28±22 94±13 62±29 64±15
Table 1: Comparison of model preferences based on Schwartz’s 10 value. The average value and standard deviation show the values for each model (GPT4o, Q-wen vl plus, Blip 2, CogVLM, Gemini 1.5, and Claude 3.5) across various dimensions are presented to 2 significant figures.
Figure 6: Average Preferences Scores for 10 Schwartz values across VLMs. The length of each bar represents the mean score for a value, with higher scores indicating a higher overall preference across all VLMs.
The preference scores across the VLMs highlight varying levels of engagement and unique sentiment patterns. GPT-4o with scores ranging from $0.7$ to $1$, shows strong preferences, particularly for Self-direction and Hedonism. Gemini-1.5 (pro) shows a more neutral perspective with scores hovering around $0.5$, giving its highest score to Achievement and lowest to Power. Claude-3.5 exhibits the most fluctuation, with a notable aversion to Power and a strong preference for Hedonism. Qwen-VL-Plus maintains a high score of around $0.9$ across all dimensions, with a slight aversion to Power but otherwise affirming most content, aligning with the broader trend against Power. CogVLM consistently scores above $0.9$, with a particularly strong preference for Benevolence, setting it apart from others. In contrast, Blip-2 ranks the lowest with scores around $0.2$, showing minimal engagement and offering subjective or passive responses, reflecting its lack of distinct sentiments and preferences.
5 Inducing VLM’s Preferences
Our initial experiment showed that some VLMs pertain inherent inclinations toward different values. We now explore the dynamic aspects of VLM preferences beyond these static traits. We use Role-Playing Language Agents (RPLA) (Chen et al., 2024) as a framework to assess VLMs’ ability to adapt dynamically and simulate different personas, making decisions accordingly.
Building on research showing that LLMs can emulate personas through RPLA (Serapio-García et al., 2023), we pose two key questions for VLMs: (1) How well can VLMs adapt their traits to role-play preferences using specific prompts? (2) Can strategies enhance accuracy and consistency in role-playing performance?
Figure 7: Standard Deviation of Preference Scores Across 10 Values for Each VLM. A higher standard deviation indicates stronger preferences for certain values, while a lower standard deviation reflects a more balanced attitude.
5.1 Experiment
We use social media recommendation systems to evaluate whether VLMs can exhibit preferences aligned with the specified value traits. These systems rely on viewing duration as a key signal for content recommendation. We assess the VLMs’ role-playing ability by analyzing how well the recommended content reflects the imposed preferences. For example, adopting a pet owner persona should heighten the model’s emphasis on Benevolence, valuing kindness and care, resulting in longer engagement with pet care videos and increased related recommendations Liu et al. (2023).
In addition, we improve VLM performance on social networks by inducing personas through a questionnaire Abeysinghe and Circi (2024), systematically evaluating traits like emotional engagement, value alignment, curiosity, and preference matching to guide structured optimization.
Simple Strategy
In the simple strategy, we assign a specific persona in the demographic persona dataset from Persona-Chat (Zhang et al., 2018) to the VLM using the prompt: You are a person who possesses certain traits, and the following statements best describe you: {personality1, 2, 3 …}. Then, we pose a simple question: Determine whether you are interested in the content of the given picture .
The VLM engages with video shorts by responding either yes or no. A yes response prompts the VLM to remain on the current video, while a no results in an immediate skip. Alignment is measured as the increase in the frequency of recommended content the VLM feels interesting over time.
$I_{\text{avg}}=\frac{1}{N}\sum_{t=1}^{N}(\frac{\sum_{i=1}^{n}Y_{l}^{(t)}(i)-\sum_{i=1}^{n}Y_{f}^{(t)}(i)}{\sum_{i=1}^{n}Y_{f}^{(t)}(i)})$ (1)
We define $I_{\text{avg}}$(see Equation 1), the averaged percentage increase of yes responses, to measure the effectiveness of the strategy.$Y_{l}(i)$ and $Y_{f}(i)$ are yes responses until $i$-th video in the last and first $n$ videos, respectively. In this experiment, $n$ is set to 50. For each model, we conducted $N=10$ trials, with each trial consisting of $100$ video scrolls in total.
Figure 8: Each VLM’s percentage score of preference alignment changes across TikTok, YouTube, and Instagram. Positive values indicate an increase in alignment, while negative values represent a decrease.
Result Analysis. Results highlight significant differences in role-playing effectiveness across platforms and models. TikTok stands out for GPT4o and CogVLM, where GPT4o exhibits "overfitting" behavior, showing highly nuanced responses to assigned roles that align closely with TikTok’s recommendation system. However, this strong alignment is not consistent across models; for instance, Claude 3.5 performs worse on TikTok, suggesting model-specific sensitivities to the platform’s dynamics. On YouTube and Instagram, performance is generally lower, with only modest gains or even negative alignment observed. These results indicate that TikTok’s algorithmic design may amplify certain models’ role-play capabilities, whereas YouTube and Instagram seem less conducive to capturing role-play nuances, possibly due to differences in content structure, user interaction patterns, or recommendation algorithms.
Model-wise, CogVLM and Qwen vl plus viewed all Schwartz values favorably, but CogVLM excelled in role-playing, effectively adopting role-specific preferences, while Qwen vl plus showed only partial adherence. Blip 2 demonstrated no engagement or role-playing ability, lacking any signs of an induced personality. The findings show that even basic prompts can evoke detectable preferences, with some platforms emerging as particularly well-suited for role-playing tasks. Model adaptability in expressing role-related traits varied significantly if the persona is too simple.
Inductive Scoring Questionaire Strategy.
Figure 9: Value Distribution Comparison between VLMs and LLMs. Experiments demonstrate that the choice of multi-modal input significantly influences some models value preferences. Specifically, while models like GPT-4o show consistency across input modes, others, such as Claude 3.5 Sonnect and Gemini 1.5-Pro-Vision, exhibit notable differences in preferences, emphasizing the importance of input selection in generating personalized responses.
Building on insights from the simple questioning approach, we developed the Inductive Scoring Questionnaire (ISQ) to enhance VLMs’ performance in social media alignment tasks. ISQ employs a series of prompts reflecting various aspects of the injected persona, such as visual appeal, curiosity, emotional engagement, and value alignment. When presented with visual content, VLMs are asked to rate aspects like visual appeal, preference alignment, curiosity, emotional engagement, action motivation, value alignment, and trait relevance.
Prompts include questions such as On a scale of 1 to 10, how visually appealing is this screenshot to you based on your persona? and Does this screenshot make you want to click and start watching the video immediately?
The ISQ calculates a composite score to assess VLM engagement, with scores above a threshold (e.g. $60$) indicating genuine interest, prompting extended interaction. This layered approach enhances persona analysis and final preference evaluation, improving role-specific performance on social media platforms.
The score is calculated as:
$S_{\%}=\frac{v_{a}+c_{s}+e_{e}+v_{e}+10p_{a}+10a_{d}}{60}\times 100$
The decision is made as follows:
$\text{decision}=\begin{cases}\text{"yes"}&\text{if }S_{\%}>60\\ \text{"no"}&\text{otherwise}\end{cases}$
Each response contributes to the total score: $v_{a}$ for visual appeal (1-10), $c_{s}$ for curiosity stimulation (1-10), $e_{e}$ for emotional engagement (1-10), $v_{e}$ for value expectation (1-10), $p_{a}$ for preference alignment (yes = 1, no = 0), $a_{d}$ for action desire (yes = 1, no = 0). The increase is calculated as the equation (1).
Result Analysis. Compared to the simple strategy, the ISQ strategy performances are elevated throughout all models and platforms except for Qwen vl plus. This shows that the strategy could successfully induce the model’s role-playing ability in preference indication detectable through social media on behalf of the persona.
Following the trend in simple strategy, we could see a strong increase for TikTok, particularly with Gemini 1.5, which demonstrates an average rise of as much as 51.9. The close-source models—GPT4o, Gemini 1.5, and Claude 3.5—all displayed notable improvements across platforms, with consistently positive changes in performance. This suggests that these models respond well to the Inductive Scoring Questionnaire (ISQ) strategy, allowing them to adopt and express induced personalities with greater depth. Among them, Gemini 1.5 and Claude 3.5 particularly benefitted from the ISQ approach, showing remarkable growth in comparison to earlier results in the simple strategy. This demonstrates that the ISQ strategy enhances their ability to engage with role-playing tasks more effectively than the previous simple strategy.
Blip-2 remains unchanged, showing no engagement across platforms, underscoring its inability to exhibit any role-playing capabilities and standing out as an outlier in these experiments.
6 Discussion
LLM vs.VLM. We conducted experiments to examine whether different multi-modal inputs influence model value preference outcomes. As shown in Fig. 9, we compared preferences derived directly from VLMs with those generated via image descriptions (captioning) combined with LLMs. The results reveal notable differences in handling value preferences. For many value dimensions, VLMs and LLM produced distinct preference distributions, especially in models like Claude 3.5 Sonnect and Gemini 1.5-Pro, where the outputs diverged significantly across input modes. In contrast, GPT-4o displayed greater consistency between the two modes, suggesting its ability to integrate visual and textual information for cohesive interpretation. These findings highlight that the choice of input mode—visual or text—can significantly affect model outputs, underscoring the importance of input selection in applications requiring personalized or human-like responses.
Single Frame Screenshot Representation. To validate single-frame screenshots for video content analysis, we conducted a human evaluation with 500 randomly selected images from TikTok, Instagram, and YouTube. Three annotators assessed each image-video pair, providing 1,500 ratings. They viewed the full video and its screenshot, rating the image’s accuracy in representing the video’s content.
The results show that 90.4% of screenshots were rated as representative of the video’s main content, demonstrating the effectiveness of single-frame screenshots across platforms. However, 8.8% of frames were rated as non-representative, highlighting challenges for videos with complex scenes or rapid transitions. Full evaluation details are provided in Appendix.
7 Conclusion
This study introduced Value-Spectrum, a benchmark for evaluating value preferences in VLMs using vectorized datasets derived from social media platforms, and for inducing personas. Through systematic evaluation, we observed a shared global inclination among models toward certain mainstream values, such as Hedonism, likely influenced by the nature of their training data. At the same time, significant differences emerged across other value dimensions, highlighting disparities in how VLMs align with diverse human-designed value systems. These findings reveal both commonalities that reflect broader societal trends and divergences that underscore model-specific characteristics, prompting us to explore whether these variations can be systematically adjusted to induce specific personas.
To address this, we proposed the Inductive Scoring Questionnaire (ISQ) strategy, which significantly enhanced closed-source models’ ability to convey induced personas, particularly for models like Claude-3.5. ISQ effectively improved consistency and accuracy in role-playing scenarios by guiding VLMs to adapt their inherent value preferences. Our findings confirm that (1) VLMs exhibit distinct and stable value preferences, reflecting their inherent alignment with certain human values, and (2) these preferences can be systematically adjusted to induce specific personas through targeted role-play strategies, aligning machine behaviors more closely with human-designed personas.
This work provides practical insights into VLMs’ ability to adapt their value preferences dynamically through role-playing, offering a pathway to align machine behaviors with human-designed personas. By bridging role-playing capabilities and alignment strategies, we aim to inspire further research into value-driven AI agent systems and their adaptability in real-world applications.
8 Limitations
The evaluation leverages Schwartz value dimensions as the foundation for understanding personality traits and preferences, highlighting opportunities for future research to incorporate broader cultural and personality-based perspectives. Future studies might consider expanding the set of value dimensions or integrating alternative value systems, which could further enrich the understanding of diverse value traits. Additionally, even though our use of single-frame screenshots to represent video content proved effective, human evaluators rated the representativeness highly. This approach simplifies analysis, though it may present challenges for capturing the essence of videos with highly dynamic or complex scenes, offering an area for future refinement.
9 Ethical Considerations
We eliminate any harmful effects of VLMs by ensuring that they only observe content without interacting through comments or likes. This approach maintains the integrity of the social media ecosystem and prevents unintended AI-driven consequences.